OpenStudy (anonymous):

What is 46% of 22? Use a fraction and express your answer in simplest form.10 3/25, 11 3/5, 8 24/25, 10 3/5

OpenStudy (mertsj):

What fraction is 46%?

OpenStudy (mertsj):

And why did you repost this question?

OpenStudy (mertsj):

What does percent mean?

OpenStudy (anonymous):

What is 46% of 22? Use a fraction and express your answer in simplest form.10 3/25, 11 3/5, 8 24/25, 10 3/5 what dont u get mert

OpenStudy (anonymous):

its a basic question

OpenStudy (anonymous):

a b c or d

OpenStudy (mertsj):

I'm trying to help you with the problem but you won't cooperate by answering my questions.

OpenStudy (anonymous):

46% with 22

OpenStudy (anonymous):

when i multiplie i get 1012

OpenStudy (mertsj):

What is 50% written as a fraction?

OpenStudy (anonymous):

1/2

OpenStudy (mertsj):

Good. Because 50% = 50/100=1/2

OpenStudy (mertsj):

Now do that for 46%

OpenStudy (anonymous):

1/3 i think?

OpenStudy (mertsj):

Try again. 46/100 and reduce it.

OpenStudy (anonymous):

ik its not 1/1 1/2 or 1/3

OpenStudy (mertsj):

Diide numerator and denominator by 2

OpenStudy (anonymous):

0.23

OpenStudy (mertsj):

\[\frac{46}{100}\div\frac{2}{2}=\frac{23}{50}\]

OpenStudy (anonymous):

oh

OpenStudy (mertsj):

And "of" in mathematics means "times" So 46% 0f 22 = \[\frac{23}{50}\times \frac{22}{1}=\]

OpenStudy (mertsj):

Now, cancel the common factor 2 out of the 22 and the 50

OpenStudy (mertsj):